# Optimal Bias-Correction and Valid Inference in High-Dimensional Ridge Regression: A Closed-Form Solution

### v1

@@ 2.4 Ridge Screening
for any $\|{\mathbf{v}}\|_{2}^{2}=c_{*}>0$ which is a finite positive constant.
Assumptions 3-4 are natural conditions about the orders of the singular values of ${\mathbf{X}}\in R^{n\times p}$ and ${\mathbf{X}}_{\mathcal{M}_{k}}$, and the penalty parameter $\lambda^{*}$ (or $\lambda$) is comparable to the magnitude of $d_{j}^{2}$ (or $d_{\mathcal{M}_{k},j}^{2}$). Similar to the illustrations for Assumption 1, the magnitude in Assumption 3-4 can be easily verified if the entries of ${\mathbf{X}}$ are independent copies of a random variable with zero mean, unit variance, and finite fourth moment under the setting $p/n\in(1,\infty)$; see Bai and Yin (1993). In fact, the orders specified in Assumptions 3-4 are only employed to establish the validity of the ridge-screening method in Theorem 5 below, and they are not the only ones capable of achieving this, so long as the rates can be properly controlled in the proof of Theorem 5. For any fixed $(p,n)$ which can be large, the efficacy of the bias-correction method and the inference methods in Section 2.5 below remain valid so long as the nonzero singular values and the penalty parameters are strictly greater than zero. Assumption 5 indicates that the minimum nonzero element in $\boldsymbol{\beta}$ cannot be too small, and the norm of the $i$-th projected coordinate $({\mathbf{U}}_{2}{\mathbf{U}}_{2}^{\prime}\boldsymbol{\beta})_{i}$ is bounded by the magnitude of its original coordinate $|\beta_{i}|$ (up to a small constant). This is actually a reasonable and intuitive assumption. For instance, in Assumption 5, if $\|\boldsymbol{\beta}\|_{2}\leq C\sqrt{s^{*}}$ (at most) because there are only $s^{*}$ nonzero elements in $\boldsymbol{\beta}$, then $\|{\mathbf{U}}_{1}{\mathbf{U}}_{1}^{\prime}\boldsymbol{\beta}\|_{2}=O_{p}(\sqrt{s^{*}})$, but it is a $p$-dimensional vector, meaning that the magnitude of each $({\mathbf{U}}_{1}{\mathbf{U}}_{1}^{\prime}\boldsymbol{\beta})_{i}$ is of order $\sqrt{s^{*}/p}$ on average. We may postulate that $\sqrt{s^{*}/p}=o(n^{-\tau})$ in such a case and Assumption 5 is even slightly weaker than this situation as we can allow that the projected coordinate to be of the same order as that of the original one. Assumption 6 is a general condition that includes Gaussian distributions as a special case.
@@
(i) Assuming the true parameter $\boldsymbol{\beta}$ belongs to the submodel $\mathcal{M}_{0}$ in (18). If $\frac{\log(p)}{n^{1-2\tau}}\rightarrow 0$, the number of iterations $k$ in the first stage satisfies $\max_{1\leq j\leq p^{*}}s^{*}(\frac{\lambda^{*}}{d_{j}^{2}+\lambda^{*}})^{k+1}\rightarrow 0$, and $n^{*}$, the number of selected elements in (20), satisfies that $n^{*}\geq s^{*}$ when $p<n$, and
@@
(i) Theorem 5 implies that the ridge-screening method is applicable in both $p<n$ and $p>n$ scenarios. Remarkably, there’s no need to specify $s^{*}$, which is the number of nonzero elements in the true $\boldsymbol{\beta}$. For $p<n$, we can simply set $p^{*}=p$ as Assumption 2 and choose $n^{*}=p$ variables which consists of all the covariates. For $p>n$, we may choose $n^{*}\asymp n$ under the assumption that $s^{*}/n\rightarrow 0$ and $p\log(p)/n^{3-2\tau}\rightarrow 0$ in an asymptotic sense, which is reasonable because $s^{*}$ is often small and we can adopt $p/n=c\in(1,\infty)$ in the setting of random matrix theory.
(ii) Theorem 5 establishes that our proposed method from Section 2.3 can completely correct the bias of the restricted ridge estimators, which is particularly helpful when $p>n$. This, in conjunction with the results of Theorem 2, confirms the versatility of our bias-correction approaches in addressing both scenarios where $p<n$ and $p>n$.
@@
It’s important to note that the convergence is guaranteed because the sequence ${\mathbf{A}}_{k}:=\sum_{j=1}^{k}\lambda^{j}({\mathbf{X}}^{\prime}{\mathbf{X}}+\lambda{\mathbf{I}}_{p})^{-j}$ converges, as demonstrated in the proof of Theorem 2 in the Appendix. Additionally, we emphasize that the number of iterations is typically not large, as only a logarithmic order of the dimension is required to ensure convergence, as discussed in Remark 2.
Finally, we briefly discuss the methods for selecting the unknown parameters in Algorithms 1-2 of the proposed approaches. Firstly, the sole unknown parameter in Algorithm 1 is $\lambda$. Throughout this paper, we assume $\lambda$ to be given, as our focus is mainly on the bias-correction issue for ridge estimators. Empirically, one can employ the widely-accepted information criteria (e.g., AIC or BIC) or utilize cross-validation methods to determine an appropriate $\lambda$. For a comprehensive understanding, readers are referred to Section 1.8 of van Wieringen (2023). Secondly, in the ridge-screening method of Algorithm 2 when $p>n$, the parameters $(\lambda^{*},n^{*})$ are unknown. Here, one can employ the aforementioned information criteria or cross-validation methods, as discussed in Sections 1.8.1-1.8.3 of van Wieringen (2023), to simultaneously determine $(\lambda^{*},n^{*})$. This can be achieved by selecting a candidate for $\lambda^{*}$ and then determining the optimal $n^{*}$ such that the pair $(\lambda^{*},n^{*})$ minimizes the information criterion or yields the best prediction performance on test sets using the restricted ridge estimator $\widehat{\boldsymbol{\beta}}_{\mathcal{M}_{k}}(\lambda^{*})$. The final choice of $(\lambda^{*},n^{*})$ can be determined by evaluating the performance across each grid point of the penalty parameters. After identifying the significant variables through the RS method, the parameter $\lambda$ used in bias-correction is equal to $\lambda^{*}$, as $\lambda^{*}$ represents the optimal choice for the associated $n^{*}$. Given that these methods are well-established in the literature, we omit further details here to save space.
@@ 2.5 Inference
In this section, we briefly introduce the inference method for the de-biased ridge estimators in Sections 2.3 and 2.4. The following theorem can be derived immediately based on the proofs of Theorem 2 and Theorem 5.
@@
(i) If $p<n$ and ${\mathbf{X}}^{\prime}{\mathbf{X}}$ is invertible. Under Assumption 1, we have
@@
(ii) Under Assumptions 2-6 and a true model in (18) when $p>n$. Suppose $k$ is sufficiently large in the ridge-screening such that the event $\{\mathcal{M}_{0}\subset\mathcal{M}_{k}(\lambda^{*})\}$ holds. Then the estimator in (22) has the following limiting distribution
@@
$\widehat{\boldsymbol{\beta}}_{c,k}(\lambda)-\boldsymbol{\beta}=O_{p}(n^{-1/2})\,\,\text{and}\,\,\widehat{\boldsymbol{\beta}}_{\mathcal{M}_{k},l}(\lambda)-\boldsymbol{\beta}_{\mathcal{M}_{k}}=O_{p}(n^{-1/2}),$
implying that our de-biased estimators are convergent with standard rate $\sqrt{n}$.
@@
$\widehat{\sigma}=\sqrt{\frac{1}{n}\|{\mathbf{y}}-{\mathbf{X}}\widehat{\boldsymbol{\beta}}_{c,k}(\lambda)\|_{2}^{2}}\,\,\text{or}\,\,\widehat{\sigma}=\sqrt{\frac{1}{n}\|{\mathbf{y}}-{\mathbf{X}}_{\mathcal{M}_{k}}\widehat{\boldsymbol{\beta}}_{\mathcal{M}_{k},l}(\lambda)\|_{2}^{2}},$
@@ 2.6 Time Series Ridge Regression
In this section, we briefly illustrate the feasibility of the proposed bias-correction method in time series ridge regression. In particular, we consider the following Auto-Regressive (AR) model:
@@
where ${\mathbf{x}}_{t}=(y_{t-1},...,y_{t-p})^{\prime}$ is the covariate vector consisting of the $p$ lagged variables of $y_{t}$, and $\boldsymbol{\beta}=(\beta_{1},...,\beta_{p})^{\prime}$ is the associated parameter vector. We restrict our consideration to cases where $p<n$. It’s uncommon to encounter situations where the number of lagged regressors in autoregressive (AR) models exceeds the sample size. In practical datasets, instances of AR models with orders surpassing, for instance, $10$, are quite rare.
@@
Equation (29) suggests that the bias-correction procedure remains applicable to time series regression with weakly dependent regressors. Additionally, Equation (30) confirms that the asymptotic distribution from Theorem 6(i) holds true for time series data as well.
2.7 Constructions of Confidence and Prediction Intervals
In this section, we explore methods for constructing confidence and prediction intervals based on ridge regression. For simplicity, we assume $\boldsymbol{\Sigma}_{\varepsilon}=\sigma^{2}{\mathbf{I}}_{n}$ to facilitate the illustrations.
We begin by outlining the construction of confidence intervals for the mean response $E(y|{\mathbf{x}}_{0})={\mathbf{x}}_{0}^{\prime}\boldsymbol{\beta}$ with a given covariate ${\mathbf{x}}_{0}$ when $p<n$. Note that $\widehat{y}_{0}={\mathbf{x}}_{0}^{\prime}\widehat{\boldsymbol{\beta}}(\lambda)$ in a ridge regression, then, by (5) in Remark 5(ii) with a sufficiently large $k$, we have
@@ 2.7 Constructions of Confidence and Prediction Intervals
The values presented in (32)-(35) are directly estimated from the data. Consequently, we are equipped to construct valid confidence and prediction intervals in ridge regression employing our proposed bias-correction techniques. Analogously, confidence and prediction intervals can also be established for models utilizing de-biased ridge estimators post ridge-screening when $p>n$. For instance, the $(1-\alpha)$-prediction interval of the future value $y_{n+1}$ produced by the restricted ridge estimator is $[L_{3},U_{3}]$ with
@@ 2.8 Bias-Variance Trade-off
The bias-variance trade-off is a fundamental concept in machine learning and statistics. It refers to the delicate balance between two sources of error in a predictive model. It is a way of analyzing a learning algorithm’s expected errors. Therefore, it would be interesting if we can derive the bias-variance trade-off of our de-biased estimators as we increase the number of iterations.
In this section, we study the MSE of the de-biased ridge estimator $\widehat{\boldsymbol{\beta}}_{c,k}(\lambda)$ in the $k$-th iteration for $p<n$ with a given $\lambda>0$. The analysis of the scenario where $p>n$ is similar, and hence, we only derive it for the case when $p<n$. The MSE of the $k$-th de-biased ridge estimator is defined as
@@
In the following theorem, we provide some sufficient conditions under which we have a bias-variance trade-off in the proposed method.
@@
Let Assumption 1 hold and ${\mathbf{U}}_{1}^{\prime}\boldsymbol{\beta}=(\delta_{1},...,\delta_{p})^{\prime}$ where $p<n$.
(i) If $\frac{\delta_{i}^{2}d_{i}^{2}}{\sigma^{2}}<1$ and $\frac{\delta_{i}^{2}d_{i}^{2}}{\sigma^{2}}+(\frac{\lambda}{\lambda+d_{i}^{2}})^{k^{*}+1}\geq 1$ for $1\leq i\leq p$ and some $k^{*}\geq 1$, then, as the number of iterations $k$ increases, the MSE of $\widehat{\boldsymbol{\beta}}_{c,k}(\lambda)$ will initially decrease to its minimum value and subsequently rise to a stable level. In particular, the minimum of the MSE can be achieved at the $k$-th iteration for some $k\in[\lfloor k_{1}\rfloor,\lfloor k_{2}\rfloor]$, where
@@
(ii) If $\frac{\delta_{i}^{2}d_{i}^{2}}{\sigma^{2}}>1$ for $1\leq i\leq p$, then, as the number of iterations $k$ increases, the MSE of $\widehat{\boldsymbol{\beta}}_{c,k}(\lambda)$ will decrease to a stable level.
@@
(i) Theorem 7 indicates there exists a bias-variance trade-off in the proposed bias-correction procedure under certain conditions. That is, if the design matrix and the penalty meet the criteria outlined in Theorem 7(i), the MSE of the de-biased estimators can attain a minimum value by balancing the bias and variance terms as the number of iterations $k$, or equivalently, the "model complexity", increases.
@@
suggesting that the proposed de-biased estimators can further minimize the MSEs while also correcting a portion of the bias compared to classical ridge estimators.

### v2

@@ 2.4 Ridge Screening
for any $\|{\mathbf{v}}\|_{2}^{2}=c_{*}>0$, which is a finite positive constant.
Assumptions 3 and 4 are natural conditions about the orders of the singular values of ${\mathbf{X}}\in R^{n\times p}$ and ${\mathbf{X}}_{\mathcal{M}_{k}}$, and the penalty parameter $\lambda^{*}$ (or $\lambda$) is comparable to the magnitude of $d_{j}^{2}$ (or $d_{\mathcal{M}_{k},j}^{2}$). Similar to the illustrations for Assumption 1, the magnitude in Assumptions 3 and 4 can be easily verified if the entries of ${\mathbf{X}}$ are independent copies of a random variable with zero mean, unit variance, and finite fourth moment under the setting $p/n\in(1,\infty)$; see Bai and Yin (1993). In fact, the orders specified in Assumptions 3 and 4 are only employed to establish the validity of the ridge-screening method in Theorem 5 below, and they are not the only ones capable of achieving this, so long as the rates can be properly controlled in the proof of Theorem 5. For any fixed $(p,n)$, which can be large, the efficacy of the bias-correction method and the inference methods in Section 2.5 below remain valid so long as the nonzero singular values and the penalty parameters are strictly greater than zero. Assumption 5 indicates that the minimum nonzero element in $\boldsymbol{\beta}$ cannot be too small, and the norm of the $i$-th projected coordinate $({\mathbf{U}}_{2}{\mathbf{U}}_{2}^{\prime}\boldsymbol{\beta})_{i}$ is bounded by the magnitude of its original coordinate $|\beta_{i}|$ (up to a small constant). This is actually a reasonable and intuitive assumption. For instance, in Assumption 5, if $\|\boldsymbol{\beta}\|_{2}\leq C\sqrt{s^{*}}$ (at most) because there are only $s^{*}$ nonzero elements in $\boldsymbol{\beta}$, then $\|{\mathbf{U}}_{1}{\mathbf{U}}_{1}^{\prime}\boldsymbol{\beta}\|_{2}=O_{p}(\sqrt{s^{*}})$, but it is a $p$-dimensional vector, meaning that the magnitude of each $({\mathbf{U}}_{1}{\mathbf{U}}_{1}^{\prime}\boldsymbol{\beta})_{i}$ is of order $\sqrt{s^{*}/p}$ on average. We may postulate that $\sqrt{s^{*}/p}=o(n^{-\tau})$ in such a case and Assumption 5 is even slightly weaker than this situation as we can allow that the projected coordinate is of the same order as that of the original one. Assumption 6 is a general condition that includes Gaussian distributions as a special case.
@@
(i) Assume the true parameter $\boldsymbol{\beta}$ belongs to the submodel $\mathcal{M}_{0}$ in (18). If $\frac{\log(p)}{n^{1-2\tau}}\rightarrow 0$, the number of iterations $k$ in the first stage satisfies $\max_{1\leq j\leq p^{*}}s^{*}(\frac{\lambda^{*}}{d_{j}^{2}+\lambda^{*}})^{k+1}\rightarrow 0$, and $n^{*}$, the number of selected elements in (20), satisfies that $n^{*}\geq s^{*}$ when $p<n$, and
@@
(i) Theorem 5 implies that the ridge-screening method is applicable in both $p<n$ and $p>n$ scenarios. Remarkably, there is no need to specify $s^{*}$, which is the number of nonzero elements in the true $\boldsymbol{\beta}$. For $p<n$, we can simply set $p^{*}=p$ as Assumption 2 and choose $n^{*}=p$ variables, which consists of all the covariates. For $p>n$, we may choose $n^{*}\asymp n$ under the assumption that $s^{*}/n\rightarrow 0$ and $p\log(p)/n^{3-2\tau}\rightarrow 0$ in an asymptotic sense, which is reasonable because $s^{*}$ is often small and we can adopt $p/n=c\in(1,\infty)$ in the setting of random matrix theory.
(ii) Theorem 5 establishes that our proposed method from Section 2.3 can completely correct the bias of the restricted ridge estimators, which is particularly helpful when $p>n$. This, in conjunction with the results of Theorem 2, confirms the versatility of our bias-correction approaches in addressing both scenarios of $p<n$ and $p>n$.
@@
It is important to note that the convergence is guaranteed because the non-random sequence ${\mathbf{A}}_{k}:=\sum_{j=1}^{k}\lambda^{j}({\mathbf{X}}^{\prime}{\mathbf{X}}+\lambda{\mathbf{I}}_{p})^{-j}$ converges, as demonstrated in the proof of Theorem 2 in the Appendix. Additionally, we emphasise that the number of iterations is typically not large, as only a logarithmic order of the dimension is required to ensure convergence, as discussed in Remark 2. Furthermore, the computation is efficient since we only need to perform the SVD on ${\mathbf{X}}^{\prime}{\mathbf{X}}+\lambda{\mathbf{I}}_{p}$ once, as we do for the classical ridge estimator. After this single SVD, the iterative process can be applied to the diagonal singular-value matrix in the expansion.
Finally, we briefly discuss the methods for selecting the unknown parameters in Algorithms 1 and 2 of the proposed approaches. Firstly, the sole unknown parameter in Algorithm 1 is $\lambda$. Throughout this paper, we assume $\lambda$ to be given, as our focus is mainly on the bias-correction issue for ridge estimators. Empirically, one can employ the widely-accepted information criteria (e.g., AIC or BIC) or utilise cross-validation methods to determine an appropriate $\lambda$. For a comprehensive understanding, readers are referred to Section 1.8 of van Wieringen (2023). Secondly, in the ridge-screening method of Algorithm 2 when $p>n$, the parameters $(\lambda^{*},n^{*})$ are unknown. Here, one can employ the aforementioned information criteria or cross-validation methods, as discussed in Sections 1.8.1-1.8.3 of van Wieringen (2023), to simultaneously determine $(\lambda^{*},n^{*})$. This can be achieved by selecting a candidate for $\lambda^{*}$ and then determining the optimal $n^{*}$ such that the pair $(\lambda^{*},n^{*})$ minimises the information criterion or yields the best prediction performance on test sets using the restricted ridge estimator $\widehat{\boldsymbol{\beta}}_{\mathcal{M}_{k}}(\lambda^{*})$. The final choice of $(\lambda^{*},n^{*})$ can be determined by evaluating the performance across each grid point of the penalty parameters. After identifying the significant variables through the RS method, the parameter $\lambda$ used in bias-correction is equal to $\lambda^{*}$, as $\lambda^{*}$ represents the optimal choice for the associated $n^{*}$. Given that these methods are well-established in the literature, we omit further details here to save space.
@@ 2.5 Inference
In this section, we briefly introduce the inference method for the de-biased ridge estimators in Sections 2.3 and 2.4. The following theorem can be derived immediately based on the proofs of Theorems 2 and 5.
@@
(i) If $p<n$ and ${\mathbf{X}}^{\prime}{\mathbf{X}}$ is invertible, then under Assumption 1, we have
@@
(ii) If $p>n$, assume that Assumptions 2-6 hold and a true model in (18). Suppose $k$ is sufficiently large in the ridge-screening such that the event $\{\mathcal{M}_{0}\subset\mathcal{M}_{k}(\lambda^{*})\}$ holds. Then the estimator in (22) has the following limiting distribution
@@
$\widehat{\boldsymbol{\beta}}_{c,k}(\lambda)-\boldsymbol{\beta}=O_{p}(n^{-1/2})\quad\text{and}\quad\widehat{\boldsymbol{\beta}}_{\mathcal{M}_{k},l}(\lambda)-\boldsymbol{\beta}_{\mathcal{M}_{k}}=O_{p}(n^{-1/2}),$
implying that our de-biased estimators are convergent with the standard rate $\sqrt{n}$.
@@
$\widehat{\sigma}=\sqrt{\frac{1}{n}\|{\mathbf{y}}-{\mathbf{X}}\widehat{\boldsymbol{\beta}}_{c,k}(\lambda)\|_{2}^{2}}\quad\text{or}\quad\widehat{\sigma}=\sqrt{\frac{1}{n}\|{\mathbf{y}}-{\mathbf{X}}_{\mathcal{M}_{k}}\widehat{\boldsymbol{\beta}}_{\mathcal{M}_{k},l}(\lambda)\|_{2}^{2}},$
@@ 2.6 Time Series Ridge Regression
In this section, we briefly illustrate the application of the proposed bias-correction method in time series ridge regression. In particular, we consider the following Autoregressive (AR) model:
@@
where ${\mathbf{x}}_{t}=(y_{t-1},...,y_{t-p})^{\prime}$ is the covariate vector consisting of the $p$ lagged variables of $y_{t}$, and $\boldsymbol{\beta}=(\beta_{1},...,\beta_{p})^{\prime}$ is the associated parameter vector. Note that $E(y_{t})=0$ for simplicity. We restrict our consideration to cases where $p<n$. It’s uncommon to encounter situations where the number of lagged regressors in an AR model exceeds the sample size. As a matter of fact, it is rare to see that the order of an AR model exceeds 20 for non-seasonal time series in empirical applications. See, for instance, Woodward et al. (2012).
@@
Equation (29) suggests that the bias-correction procedure remains applicable to time series regression with weakly dependent regressors. Additionally, Equation (30) confirms that the asymptotic distribution of Theorem 6(i) continues to hold for time series data as well.
2.7 Construction of Confidence and Prediction Intervals
In this section, we explore methods for constructing confidence and prediction intervals based on ridge regression. For simplicity, we assume $\boldsymbol{\Sigma}_{\varepsilon}=\sigma^{2}{\mathbf{I}}_{n}$ to facilitate the illustration.
We begin by outlining the construction of confidence intervals for the mean response $E(y|{\mathbf{x}}_{0})={\mathbf{x}}_{0}^{\prime}\boldsymbol{\beta}$ with a given covariate ${\mathbf{x}}_{0}$ when $p<n$. Note that $\widehat{y}_{0}={\mathbf{x}}_{0}^{\prime}\widehat{\boldsymbol{\beta}}(\lambda)$ on a ridge regression, then, by (5) in Remark 5(ii) with a sufficiently large $k$, we have
@@ 2.7 Construction of Confidence and Prediction Intervals
The values presented in (32)-(35) are directly estimated from the data. Consequently, we are equipped to construct valid confidence and prediction intervals in ridge regression via employing the proposed bias-correction techniques. Analogously, confidence and prediction intervals can also be established for models utilising de-biased ridge estimators post ridge-screening when $p>n$. For instance, the $(1-\alpha)$-prediction interval of the future value $y_{n+1}$ produced by the restricted ridge estimator is $[L_{3},U_{3}]$ with
@@ 2.8 Bias-Variance Trade-off
The bias-variance trade-off is a fundamental concept in machine learning and statistics. It refers to the delicate balance between two sources of error in a predictive model. It is a way of analysing a learning algorithm’s expected errors. Therefore, it would be interesting if we can derive the bias-variance trade-off of our de-biased estimators as we increase the number of iterations.
In this section, we study the MSE of the de-biased ridge estimator $\widehat{\boldsymbol{\beta}}_{c,k}(\lambda)$ in the $k$-th iteration for $p<n$ with a given $\lambda>0$. The analysis of the scenario with $p>n$ is similar, and hence, we only focus on the case of $p<n$. The MSE of the $k$-th de-biased ridge estimator is defined as
@@
In the following theorem, we provide some sufficient conditions under which we have a bias-variance trade-off for the proposed method.
@@
Let Assumption 1 hold and ${\mathbf{U}}_{1}^{\prime}\boldsymbol{\beta}=(\delta_{1},...,\delta_{p})^{\prime}$, where $p<n$.
(i) If $\frac{\delta_{i}^{2}d_{i}^{2}}{\sigma^{2}}<1$ and $\frac{\delta_{i}^{2}d_{i}^{2}}{\sigma^{2}}+(\frac{\lambda}{\lambda+d_{i}^{2}})^{k^{*}+1}\geq 1$, for $1\leq i\leq p$, and some $k^{*}\geq 1$, then, as the number of iterations $k$ increases, the MSE of $\widehat{\boldsymbol{\beta}}_{c,k}(\lambda)$ will initially decrease to its minimum value and subsequently rise to a stable level. In particular, the minimum of the MSE can be achieved at the $k$-th iteration for some $k\in[\lfloor k_{1}\rfloor,\lfloor k_{2}\rfloor]$, where
@@
(ii) If $\frac{\delta_{i}^{2}d_{i}^{2}}{\sigma^{2}}>1$, for $1\leq i\leq p$, then, as the number of iterations $k$ increases, the MSE of $\widehat{\boldsymbol{\beta}}_{c,k}(\lambda)$ will decrease to a stable level.
@@
(i) Theorem 7 indicates there exists a bias-variance trade-off in the proposed bias-correction procedure under certain conditions. That is, if the design matrix and the penalty satisfy the criteria outlined in Theorem 7(i), the MSE of the de-biased estimators can attain a minimum value by balancing the bias and variance terms as the number of iterations $k$, or equivalently, the "model complexity", increases.
@@
suggesting that the proposed de-biased estimators can further minimise the MSEs while also correct a portion of the bias compared to classical ridge estimators.